OpenStudy (anonymous):

The graph below represents which system of inequalities? graph of two intersecting lines. One line is solid, and goes through the points (0, negative 2) and (1, 0) and is shaded in below the line. The other line is dashed, and goes through the points (0, 3) (3, 0) and is shaded in above the line. y > 2x - 3 y > -x - 3 y < 2x - 2 y < -x + 3 y less than or greater to 2x - 2 y > -x + 3 None of the above

OpenStudy (anonymous):

@mtbender74 @Peabody

OpenStudy (anonymous):

@ganeshie8 @AccessDenied @iPwnBunnies @Hero plz help

ganeshie8 (ganeshie8):

First get the equations for corresponding lines in `slope-intercept` form

ganeshie8 (ganeshie8):

The solid line rising up by 2 units for every 1 unit to right, so the slope is `2` Also it is cutting the y axis at -2, that makes the y-intercept = `-2` So the equation for solid line is : `y = 2x - 2`

ganeshie8 (ganeshie8):

Can you try writing out the equation for dotted line ?

OpenStudy (anonymous):

Yes ill try

ganeshie8 (ganeshie8):

good, get the slope and y-intercept first

OpenStudy (anonymous):

The slope for the dotted line is -1 because it is decreasing by one and its rise/run

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

How do you get the y intercept again I forgot...

ganeshie8 (ganeshie8):

"intercept" is a fancy word for "cut"

ganeshie8 (ganeshie8):

its the place at which the line cuts the y axis

OpenStudy (anonymous):

3?

ganeshie8 (ganeshie8):

right !!

OpenStudy (anonymous):

Y=Mx + B

OpenStudy (anonymous):

Y=-1x + 3

ganeshie8 (ganeshie8):

Looks good !

ganeshie8 (ganeshie8):

so the corresponding equations are : y = 2x - 2 y = -x + 3 right ?

OpenStudy (anonymous):

wait how did you get the first equation?

ganeshie8 (ganeshie8):

we're almost done ! just need to fix the inequality based on the shading shown in the graph

ganeshie8 (ganeshie8):

look at my second reply from top

OpenStudy (anonymous):

Oh ok XD didn't see that

OpenStudy (anonymous):

So its C?

ganeshie8 (ganeshie8):

lol okay :) knw how to fix the inequality based on shading ?

OpenStudy (anonymous):

No not really I kinda just go on a graphing app and get the shading from there

ganeshie8 (ganeshie8):

just put `<` symbol if the shading is `DOWN`

OpenStudy (anonymous):

Yes but there are different shadings what do those mean

ganeshie8 (ganeshie8):

look at the graph again, the shading for solid line is `DOWN` the line - so the inequality is : \( \large y \le 2x-2\)

ganeshie8 (ganeshie8):

Whereas the shading for dotted line is `ABOVE` the line - so the inequality is : \(\large y > -x + 3\)

OpenStudy (anonymous):

Oh ok I got it I think

ganeshie8 (ganeshie8):

great ! just be careful on when to use \(\le \) and just \(\lt\)

OpenStudy (anonymous):

How do you get that sign on the keboard I know how to get <> but I ont know the other one

ganeshie8 (ganeshie8):

here it is : solid line : \(\le, \ge\) dotted line : \(\lt, \gt\)

ganeshie8 (ganeshie8):

oh you need to use latex code

ganeshie8 (ganeshie8):

copy paste exact below in ur next reply : ``` \( \le \) ```

OpenStudy (anonymous):

\(

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

copy paste the entire thingy

OpenStudy (anonymous):

oh ok ike this? \( \le \)

OpenStudy (anonymous):

it worked!

OpenStudy (anonymous):

Thank You! and just to make sure its c right?

ganeshie8 (ganeshie8):

Congrats ! yes C is right :)

OpenStudy (anonymous):

Thanks

ganeshie8 (ganeshie8):

yw